# On De Giorgi’s lemma for variational interpolants

### v2

@@ 3.1 Setup
We now collect our working assumptions on ${\mathcal{E}}$ and ${\mathcal{R}}$, which partly mirror those collected in Definition 2.3 and Hypothesis 2.5. At the same time, they clearly reflect the underlying Banach setup, involving the weak topology on $X$, in addition to the norm topology, in conditions (3.3) below (cf. also Remark 2.6). As already mentioned in the Introduction, in the Banach setting we will allow for nonsmooth energies, and thus work with the Fréchet subdiffential $\partial{\mathcal{E}}$ of ${\mathcal{E}}$ in place of its Gâteau derivative ${\mathrm{D}}{\mathcal{E}}$. We recall that the multivalued operator $\partial{\mathcal{E}}:X\rightrightarrows X^{*}$ is defined at $u\in\mathrm{dom}({\mathcal{E}})$ by
@@
Hypothesis 3.1 (Conditions for generalized Banach-space gradient systems)
@@
  • The dissipation potential ${\mathcal{R}}:X\to[0,\infty)$ is lower semicontinuous, convex, and fulfills, along with its convex conjugate ${\mathcal{R}}^{*}:X^{*}\to[0,+\infty)$, the following conditions:
    ${\mathcal{R}}(0)=0\,,$ ${\mathcal{R}}^{*}(0)=0\,,$ (3.2)
    $\lim_{\|v\|\to\infty}\frac{{\mathcal{R}}(v)}{\|v\|}=\infty,$ $\lim_{\|\omega\|_{*}\to\infty}\frac{{\mathcal{R}}^{*}(\omega)}{\|\omega\|_{*}}=\infty.$
@@
It is often significant to consider dissipation potentials that also depend on the state variable, i.e. ${\mathcal{R}}={\mathcal{R}}(u,v)$, but this generalization would be irrelevant for the study of the properties of the single-step minimum problem
$\mathrm{Min}_{u\in X}\Phi_{\sigma}(u^{o};u)\qquad\text{with }\ \Phi_{\sigma}(u^{o};u):=\sigma{\mathcal{R}}\left(\frac{1}{\sigma}(u{-}u^{o})\right)+{\mathcal{E}}(u),\quad\sigma>0,$ (3.4)
for a fixed $u^{o}\in M$. Indeed, in the state-dependent case the dissipation term would be simply replaced by ${\mathcal{R}}\left(u^{o},\frac{1}{\sigma}(u{-}u^{o})\right)$.
@@
We will stick to notation (2.5), namely
  • $\phi=\phi(u^{o};\sigma)$ for the value functional associated with the above minimum problem; in fact, we shall also refer to $\phi(u^{o};\cdot)$ as marginal function. We remark for later use that, in analogy to the metric case in [AGS05], it was proved in [MRS13b, Lem. 6.1] that
    $\lim_{\sigma\downarrow 0}\phi(u^{o};\sigma)={\mathcal{E}}(u^{o})\,.$ (3.5)
  • and $J_{\sigma}(u^{o})$ for the set of minimizers.
Additionally, in the Banach setup under Hypothesis 3.1 we have at out disposal that every $\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ satisfies the Euler-Lagrange equation for (3.4), namely
$0\in\partial{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\widetilde{\xi}_{\sigma},\quad\widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma}).$ (3.6)
Furthermore, observe that, under Hypothesis 3.1 the following holds:
@@
    We have $J_{\sigma}(u^{o})\neq\emptyset$ for all $u^{o}\in\mathrm{dom}({\mathcal{E}})$ and all $\sigma>0$.
    For this, it suffices to observe that any infimizing sequence for $\Phi_{\sigma}(u^{o};\cdot)$ is bounded in $X$ (thanks to the facts that ${\mathcal{E}}$ is bounded from below and ${\mathcal{R}}$ has superlinear growth), and to resort to the weak lower semicontinuity of ${\mathcal{E}}$ and ${\mathcal{R}}$.
@@
    The multi-valued mapping $\Gamma\colon(0,\infty)\rightrightarrows X$; $\Gamma(\sigma):=J_{\sigma}(u^{o})$ is upper semicontinuous from ${\mathbb{R}}$ to $X$, in the sense that inclusion (2.9a) holds (even for the Kuratowski upper limit $\mathop{\mathrm{Ls}}_{n\to\infty}^{\mathrm{weak}}\Gamma(\sigma_{n})$ defined in terms of the weak topology on $X$). Therefore, by [CaV77, Thm. 3.22] we may conclude the existence of a (strongly) measurable selection $(0,\infty)\ni\sigma\to\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$.
@@
We consider
  • the ambient space $X={\mathrm{L}}^{p}(\Omega)$, $p>1$, with a bounded Lipschitz domain $\Omega\subset{\mathbb{R}}^{d}$;
  • the dissipation potential ${\mathcal{R}}:{\mathrm{L}}^{p}(\Omega)\to[0,\infty)$ defined by ${\mathcal{R}}(v)=\int_{\Omega}{\mathrm{R}}(v(x))\;\!\mathrm{d}x$ with
@@
3.2 De Giorgi’s estimate and identity in the Banach setup
In order to state the Banach-space versions of the metric energy identity (2.11) and of De Giorgi’s estimate (2.13), we are naturally led to introduce the following object, which corresponds to the “slope part of the dissipation” in (2.13).
@@ 3.2 De Giorgi’s estimate and identity in the Banach setup
Let $u\in\mathrm{dom}(\partial{\mathcal{E}})$. The quantity
@@
Indeed, as a consequence of the superlinear growth of ${\mathcal{R}}^{*}$, guaranteeing that infimizing sequences for the above minimum problem are bounded in $X^{*}$, and of the closedness property (3.3d), the infimum in (3.9) is attained, i.e. we have that
@@
Likewise, it is immediate to check that, again by (3.3d), for every $(u_{n})_{n},\,u\in X$
$u_{n}\rightharpoonup u\ \ \Longrightarrow\ \ \liminf_{n\to\infty}\mathcal{S}_{\mathcal{R}}(u_{n})\geq\mathcal{S}_{\mathcal{R}}(u)\,.$ (3.10b)
Obviously, one may expect that the ${\mathcal{R}}$-slope $\mathcal{S}_{\mathcal{R}}(\widetilde{u}_{\sigma})$, evaluated along a (measurable) selection $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$, will play the role of the term $\psi^{*}\big{(}\big{|}\partial{\mathcal{E}}\big{|}(\widetilde{u}_{\sigma})\big{)}$ in the Banach-space version of (2.13). However, a more careful comparison with the metric identity (2.11), featuring the quantity
$\tfrac{1}{\sigma}\mathsf{d}_{\sigma}^{-}(u^{o})=\inf\big{\{}\,\|\tfrac{1}{\sigma}(u{-}u^{o})\|\>\big{|}\>u\in J_{\sigma}(u^{o})\,\big{\}},$
suggests that the notion of slope has to be adjusted. In fact, it is expedient to bring into the picture the additional structure available in the Banach setup, namely the fact that every $u\in J_{\sigma}(u^{o})$ fulfills the Euler-Lagrange equation (3.6). We thus introduce a “conditioned slope part of the dissipation”, defined by the minimization of the dual dissipation potential ${\mathcal{R}}^{*}$ over selections $\xi\in\partial{\mathcal{E}}(u)$ that additionally satisfy the Euler-Lagrange equation. Accordingly, we define a multi-valued operator which encodes the validity of (3.6). For brevity, we shall refer to these two objects as conditioned ${\mathcal{R}}$-slope and conditioned subdifferential, respectively.
@@
The multivalued mapping $\partial_{\mathcal{R}}{\mathcal{E}}:X{\times}(0,\infty){\times}X\rightrightarrows X^{*}$ defined by
$\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u):=\Big{\{}\>\xi\in X^{*}\;\Big{|}\;\xi\in\partial{\mathcal{E}}(u)\text{ and }{-}\xi\in\partial{\mathcal{R}}\big{(}\frac{1}{\sigma}(u{-}u^{o})\big{)}\>\Big{\}}$ (3.11)
@@
$\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u):=\inf\big{\{}\,{\mathcal{R}}^{*}({-}\xi)\>\big{|}\>\xi\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\,\big{\}}$ (3.12)
is called conditioned ${\mathcal{R}}$-slope of ${\mathcal{E}}$ at $u\in\mathrm{dom}(\partial{\mathcal{E}})$.
@@
Clearly, since every $u\in J_{\sigma}(u^{o})$ fulfills the Euler-Lagrange equation (3.6), we have $\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;u)\neq\emptyset$ and thus $\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u)<\infty$. Obviously, in general there holds
$\mathcal{C}_{\mathcal{R}}(u^{o},\sigma;u)\geq\mathcal{S}_{\mathcal{R}}(u)\qquad\text{for all }u\in J_{\sigma}(u^{o}),$
which seems to suggest $\mathcal{C}_{\mathcal{R}}$ as the ‘right’ object for the attainment of an equality in De Giorgi’s estimate, cf. also Example 3.6 below.
@@
We say that a (strongly) measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}(u^{o})$ fulfills, on some interval $(0,\sigma_{*})$,
  • De Giorgi’s estimate
    if there holds for every $\sigma\in(0,\sigma_{*})$
    ${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{\rho=0}^{\sigma}\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{o});$ (3.13)
  • De Giorgi’s identity
    if (3.13) holds as an equality, i.e. for every $\sigma\in(0,\sigma_{*})$
    ${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{\rho=0}^{\sigma}\mathcal{C}_{\mathcal{R}}(u^{o},\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{o})\,.$ (3.14)
Later on, in Section 4.3 we will provide a sufficient condition for the attainment of the $\inf$ in the definition (3.12) of $\mathcal{C}_{\mathcal{R}}$, which is also related to the validity of the analog of the lower semicontinuity properties (3.10b). We will also discuss the existence of measurable selections
$(0,\infty)\ni\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{\mathcal{R}}{\mathcal{E}}(u^{o},\sigma;\widetilde{u}_{\sigma}),\quad\text{with $\sigma\mapsto\widetilde{u}_{\sigma}$ a measurable selection in $J_{\sigma}(u^{o})$,}$
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big{(}\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}+\int_{0}^{\sigma}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{o})\,,$
@@
We conclude this section by using Example 2.13, revisited in the Banach setup, to convey the idea that the finer information encoded in the conditioned ${\mathcal{R}}$-slope of the energy may play a key role for the attainment of De Giorgi’s identity.
@@
We now treat Example (2.13) as a $\mathrm{gBGS}$ in the Hilbert or Banach space $X={\mathbb{R}}$:
$M={\mathbb{R}},\quad{\mathcal{E}}(u)=\max\{u,0\}\doteq u^{+},\quad{\mathcal{R}}(v)=\frac{1}{2}v^{2}.$
Since ${\mathcal{E}}$ is convex, its Fréchet subdifferential $\partial{\mathcal{E}}:{\mathbb{R}}\rightrightarrows{\mathbb{R}}$ coincides with that in the sense of convex analysis, and it is set-valued with $\partial{\mathcal{E}}(0)=[0,1]$. The ${\mathcal{R}}$-slope of the energy is given by
@@
As before, starting from $u^{o}=1$ we calculate the variarional interpolant $\widetilde{u}_{\sigma}=(1{-}\sigma)^{+}=\max\{1{-}\sigma,0\}$, so that
@@
${\mathcal{E}}(u^{o})-{\mathcal{E}}(\widetilde{u}_{\sigma})-\sigma{\mathcal{R}}\big{(}\tfrac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{o})\big{)}=1{-}(1{-}\sigma)^{+}-\tfrac{1}{2\sigma}|(1{-}\sigma)^{+}{-}1|^{2}=\begin{cases}\frac{1}{2}\sigma&\text{if }\sigma\leq 1,\\ 1{-}\frac{1}{2\sigma}&\text{if }\sigma\geq 1.\end{cases}$

### v3

@@ 3.1 Setup
We will thus study the minimum problem
$\mathop{\mathrm{Min}}\limits_{u\in X}\Phi_{\sigma}(u)\qquad\text{with }\ \Phi_{\sigma}(u):=\sigma{\mathcal{R}}\Big(\frac{1}{\sigma}(u{-}u^{\circ})\Big)+{\mathcal{E}}(u),\quad\sigma>0,$ (3.1)
for a fixed $u^{\circ}\in\mathop{\mathrm{dom}}({\mathcal{E}})\subset X$.
We now collect our working assumptions on $\mathcal{E}$} and $\mathcal{R}$}, which partly mirror those collected in Definition 2.3 and Hypothesis 2.5. At the same time, they clearly reflect the underlying Banach setup, involving the weak topology on $X$, in addition to the norm topology, in conditions (3.4) below (cf. also Remark 2.6). As already mentioned in the Introduction, in the Banach setting we will allow for nonsmooth energies, and thus work with the Fréchet subdifferential $\partial{\mathcal{E}}$ of $\mathcal{E}$} in place of its Gâteau derivative $\mathrm{D}$$\mathcal{E}$}. We recall that the multivalued operator $\partial{\mathcal{E}}:X\rightrightarrows X^{*}$ is defined at $u\in\mathop{\mathrm{dom}}({\mathcal{E}})$ by
@@
Hypothesis 3.1 (Conditions for $\mathrm{gBGS}$)
@@
  • The dissipation potential ${\mathcal{R}}:X\to[0,\infty]$ has a proper domain $\mathop{\mathrm{dom}}({\mathcal{R}})$ open in $X$, it is lower semicontinuous, convex, and fulfills the following conditions:
    ${\mathcal{R}}(0)=0,\qquad\lim_{\|v\|\to\infty}\frac{{\mathcal{R}}(v)}{\|v\|}=\infty.$ (3.3)
@@
We pin down some elementary, but important, consequences of our assumptions on $\mathcal{R}$}, for its convex conjugate
${\mathcal{R}}^{*}:X^{*}\to[0,\infty],\qquad{\mathcal{R}}^{*}(\xi)=\sup_{v\in X}(\langle\xi,v\rangle{-}{\mathcal{R}}(v)),$
which is, by construction, lower semicontinuous and convex on $X^{*}$.
Lemma 3.2
The functional ${\mathcal{R}}^{*}:X^{*}\to[0,\infty]$ has domain $\mathop{\mathrm{dom}}({\mathcal{R}}^{*})=X^{*}$, it fulfills ${\mathcal{R}}^{*}(0)=0$, and the coercivity estimate
$\exists\,r,K_{r}>0\ \forall\,\xi\in X^{*}\,:\quad{\mathcal{R}}^{*}(\xi)\geq r\|\xi\|_{*}-K_{r}\,.$ (3.5)
Furthermore, $\mathcal{R}$} is locally Lipschitz on $\mathop{\mathrm{dom}}({\mathcal{R}})$.
Proof. Since $\mathcal{R}$} has superlinear growth at infinity, $\mathcal{R}$$^{*}$ has full domain (cf., e.g., [Bre73, Prop. 2.11, 2.14]). It can be immediately checked that ${\mathcal{R}}^{*}(0)=0$. To show (3.5), we observe that since 0 belongs to the open set $\mathop{\mathrm{dom}}({\mathcal{R}})$, $\mathcal{R}$} is continuous in 0 [EkT99, Chap. I, Cor. 2.5] and, thus, there exists $r>0$ such that $\overline{B_{r}(0)}\subset\mathop{\mathrm{dom}}({\mathcal{R}})$ and $\mathcal{R}$} is bounded from above on $\overline{B_{r}(0)}$. Therefore,
${\mathcal{R}}^{*}(\xi)\geq\langle\xi,\ell rw\rangle-{\mathcal{R}}(\ell rw)\qquad\text{for all }w\in\overline{B_{1}(0)},\ \ell\in[0,1]\,.$
Taking into account that ${\mathcal{R}}(\ell rw)\leq\ell{\mathcal{R}}(rw)\leq{\mathcal{R}}(rw)$, we thus conclude
$r\|\xi\|_{*}=r\sup_{w\in\overline{B_{1}(0)}}|\langle\xi,w\rangle|\leq{\mathcal{R}}^{*}(\xi)+K\qquad\text{with }K_{r}=\sup_{v\in\overline{B_{r}(0)}}{\mathcal{R}}(v)<\infty\,.$
The last assertion follows from [EkT99, Chap. I, Cor. 2.4], which states that $\mathcal{R}$} is locally Lipschitz on (the open) $\mathop{\mathrm{dom}}({\mathcal{R}})$ if and only if there exists a non-empty convex set over which $\mathcal{R}$} is bounded from above which, in our case, is the aforementioned ball $B_{r}(0)$.
Remark 3.3
For most part of the subsequent analysis, the open-domain condition for $\mathcal{R}$} will be sufficient. One the one hand it implies the coercivity property (3.5) for $\mathcal{R}$$^{*}$, and on the other hand it implies local Lipschitz continuity of $\mathcal{R}$} in all points of the domain. Only for specific results (see Lemma 4.7 and Proposition 4.14 ahead), we will have to require that $\mathcal{R}$$^{*}$ has superlinear growth at infinity, which in particular implies that $\mathcal{R}$} has full domain.
It remains an open problem, though, to deal with the case in which $\mathop{\mathrm{dom}}({\mathcal{R}})$ is a general proper subset of $X$, not necessarily open, which occurs, for instance, when the gradient system $(X,{\mathcal{E}},{\mathcal{R}})$ models the unidirectional evolution of inelastic processes in solid mechanics like damage.
Remark 3.4
It is often significant to consider dissipation potentials that also depend on the state variable, i.e. ${\mathcal{R}}={\mathcal{R}}(u,v)$, but this generalization would be irrelevant for the study of the properties of the single-step minimum problem (3.1) for a fixed $u^{\circ}\in X$. Indeed, in the state-dependent case the dissipation term would be simply replaced by ${\mathcal{R}}\left(u^{\circ},\frac{1}{\sigma}(u{-}u^{\circ})\right)$.
@@
Nonetheless, it would also be possible to bypass (3.4d) by working with the so-called limiting subdifferential of $\mathcal{E}$}, defined at a given $u\in\mathop{\mathrm{dom}}({\mathcal{E}})$ as the set of all $\xi$ that are weak limits of sequences $(\xi_{n})_{n}$, with $\xi_{n}\in\partial{\mathcal{E}}(u_{n})$ for every $n\in{\mathbb{N}}$, and $u_{n}\to u$ with $\sup_{n}{\mathcal{E}}(u_{n})<\infty$. Gradient flows and generalized gradient systems featuring this subdifferential notion were analyzed in [RoSa06] and [MRS13], respectively.
For the minimum problem (3.1), we will stick to notation (2.5), namely
  • $\phi=\phi(\sigma)$ for the value functional associated with the above minimum problem; in fact, we shall also refer to $\phi(u^{\circ};\cdot)$ as marginal function. We remark for later use that, in analogy to the metric case in [AGS05], it was proved in [MRS13, Lem. 6.1] that
    $\lim_{\sigma\downarrow 0}\phi(\sigma)={\mathcal{E}}(u^{\circ})\,$, (3.6)
    where the superlinearity of $\mathcal{R}$} in (3.3) is essential.
  • and $J_{\sigma}$ for the set of minimizers.
Additionally, in the Banach setup under Hypothesis 3.1 every $\widetilde{u}_{\sigma}\in J_{\sigma}$ satisfies the Euler-Lagrange equation for (3.1), namely
$0\in\partial{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\widetilde{\xi}_{\sigma}\quad\text{ with }\ \widetilde{\xi}_{\sigma}\in\partial{\mathcal{E}}(\widetilde{u}_{\sigma}).$ (3.7)
This follows as in [MRS13, Prop. 4.2] (which, in turn, relies on results from [Mor06]), whose proof also uses that $\mathcal{R}$} is locally Lipschitz around the point $\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})$; note that this is granted by Lemma 3.2. In fact, [MRS13, Prop. 4.2] would apply also if we worked with the notion of limiting subdifferential previously mentioned.
Furthermore, under Hypothesis 3.1 the following holds:
@@
    We have $J_{\sigma}\neq\emptyset$ for all $u^{\circ}\in\mathop{\mathrm{dom}}({\mathcal{E}})$ and all $\sigma>0$.
    For this, it suffices to observe that any infimizing sequence for $\Phi_{\sigma}(\cdot)$ is bounded in $X$ (thanks to the facts that $\mathcal{E}$} is bounded from below and $\mathcal{R}$} has superlinear growth), and to resort to the weak lower semicontinuity of $\mathcal{E}$} and $\mathcal{R}$}.
@@
    The multivalued mapping $\bm{J}\colon(0,\infty)\rightrightarrows X$; $\sigma\mapsto\bm{J}(\sigma):=J_{\sigma}$ (cf. (2.9)) is upper semicontinuous from $\mathbb{R}$} to $X$, in the sense that inclusion (2.10a) holds (even for the Kuratowski upper limit $\mathop{\mathrm{Ls}}_{n\to\infty}^{\mathrm{weak}}\bm{J}(\sigma_{n})$ defined in terms of the weak topology on $X$). Therefore, by [CaV77, Thm. 3.22] we may conclude the existence of a (strongly) measurable selection $(0,\infty)\ni\sigma\to\widetilde{u}_{\sigma}\in J_{\sigma}$.
@@
We consider a $\mathrm{gBGS}$ $(X,{\mathcal{E}},{\mathcal{R}})$ such that
  • $X={\mathrm{L}}^{p}(\Omega)$, $p>1$, with a bounded Lipschitz domain $\Omega\subset{\mathbb{R}}^{d}$;
  • the dissipation potential ${\mathcal{R}}:{\mathrm{L}}^{p}(\Omega)\to[0,\infty)$ is defined by ${\mathcal{R}}(v)=\int_{\Omega}{\mathrm{R}}(v(x))\;\!\mathrm{d}x$ with
@@
3.2 The De Giorgi estimate and identity in the Banach setup
In order to state the Banach-space versions of the metric energy identity (2.12) and of the De Giorgi estimate (2.14), we are naturally led to introduce the following object, which corresponds to the “slope part of the dissipation” in (2.14).
@@ 3.2 The De Giorgi estimate and identity in the Banach setup
Let $u\in\mathop{\mathrm{dom}}(\partial{\mathcal{E}})$. The quantity
@@
Indeed, as a consequence of the coercivity property (3.5) of $\mathcal{R}$$^{*}$, guaranteeing that infimizing sequences for the above minimum problem are bounded in $X^{*}$, and of the closedness property (3.4d), the infimum in (3.10) is attained, i.e. we have that
@@
To see this, consider $\xi_{n}\in\partial{\mathcal{E}}(u)$ with ${\mathcal{R}}^{*}(-\xi_{n})\to\mathcal{S}_{{\mathcal{R}}}(u)$, then the coercivity of $\mathcal{R}$$^{*}$ allows to extract a subsequence (not relabeled) such that $\xi_{n}\rightharpoonup\xi$ and the sequence $(u,\xi_{n})\rightharpoonup(u,\xi)$ satisfies the assumptions of (3.4d) with energy level $E={\mathcal{E}}(u)$. Hence, we have $\xi\in\partial{\mathcal{E}}(u)$ and $\mathcal{S}_{{\mathcal{R}}}(u)={\mathcal{R}}^{*}(-\xi)$ follows. Likewise, it is immediate to check that, for every $(u_{n})_{n},\,u\in X$ in some energy sublevel (recall that the closedness property (3.4d) holds under an energy bound), we have
$(u_{n})_{n},\,u\in S_{E},\ \ u_{n}\rightharpoonup u\ \ \Longrightarrow\ \ \liminf_{n\to\infty}\mathcal{S}_{{\mathcal{R}}}(u_{n})\geq\mathcal{S}_{{\mathcal{R}}}(u)\,.$ (3.11b)
Obviously, one may expect that the $\mathcal{R}$}-slope $\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\sigma})$, evaluated along a (measurable) selection $\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$, will play the role of the term $\psi^{*}\big(\big|\partial{\mathcal{E}}\big|(\widetilde{u}_{\sigma})\big)$ in the Banach-space version of (2.14). However, a more careful comparison with the metric identity (2.12), featuring the quantity
$\tfrac{1}{\sigma}\mathsf{d}_{\sigma}^{-}=\inf\big\{\,\|\tfrac{1}{\sigma}(u{-}u^{\circ})\|\>\big|\>u\in J_{\sigma}\,\big\},$
suggests that the notion of slope has to be adjusted. In fact, it is expedient to bring into the picture the additional structure available in the Banach setup, namely the fact that every $u\in J_{\sigma}$ fulfills the Euler-Lagrange equation (3.7).
We thus set forth a “conditioned slope part of the dissipation”, defined by the minimization of the dual dissipation potential $\mathcal{R}$$^{*}$ over selections $\xi\in\partial{\mathcal{E}}(u)$ that additionally satisfy the Euler-Lagrange equation. Accordingly, we introduce a multivalued operator which encodes the validity of (3.7). We shall refer to these two objects as conditioned $\mathcal{R}$}-slope and conditioned subdifferential, respectively. Notice that the conditioned $\mathcal{R}$}-slope is naturally defined on the graph of the multivalued operator $\bm{J}$, namely on the set
$\mathrm{Graph}(\bm{J})=\{(\sigma,u)\in(0,\infty){\times}X\,:\ u\in\bm{J}(\sigma)=J_{\sigma}\}\,.$
@@
The multivalued mapping $\partial_{{\mathcal{R}}}{\mathcal{E}}:\mathrm{Graph}(\bm{J})\rightrightarrows X^{*}$ defined by
$\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u):=\Big\{\>\xi\in X^{*}\;\Big|\;\xi\in\partial{\mathcal{E}}(u)\text{ and }{-}\xi\in\partial{\mathcal{R}}\big(\frac{1}{\sigma}(u{-}u^{\circ})\big)\>\Big\}$ (3.12)
@@
$\mathcal{C}_{{\mathcal{R}}}(\sigma;u):=\inf\big\{\,{\mathcal{R}}^{*}({-}\xi)\>\big|\>\xi\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)\,\big\}\quad\text{for all }(\sigma,u)\in\mathrm{Graph}(\bm{J})$ (3.13)
is called conditioned $\mathcal{R}$}-slope of $\mathcal{E}$} at $u\in\mathop{\mathrm{dom}}(\partial{\mathcal{E}})$.
@@
Clearly, since every $u\in J_{\sigma}$ fulfills the Euler-Lagrange equation (3.7), we have $\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;u)\neq\emptyset$ and thus $\mathcal{C}_{{\mathcal{R}}}(\sigma;u)<\infty$ for all $(\sigma,u)\in\mathrm{Graph}(\bm{J})$. Obviously, in general there holds
$\mathcal{C}_{{\mathcal{R}}}(\sigma;u)\geq\mathcal{S}_{{\mathcal{R}}}(u)\qquad\text{for all }(\sigma,u)\in\mathrm{Graph}(\bm{J}),$ (3.14)
which seems to suggest $\mathcal{C}_{{\mathcal{R}}}$ as the ‘right’ object for the attainment of an equality in the De Giorgi estimate, cf. also Example 3.9 below.
@@
We say that a (strongly) measurable selection $(0,\infty)\ni\sigma\mapsto\widetilde{u}_{\sigma}\in J_{\sigma}$ fulfills, on some interval $(0,\sigma_{*})$,
  • the simple De Giorgi estimate
    if for all $\sigma\in(0,\sigma_{*})$ we have
    ${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{\rho=0}^{\sigma}\mathcal{S}_{{\mathcal{R}}}(\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ});$} (3.15)
  • the improved De Giorgi estimate
    if for all $\sigma\in(0,\sigma_{*})$ we have
    ${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{\rho=0}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ});$ (3.16)
  • the De Giorgi identity
    if (3.16) holds as an equality, i.e. for all $\sigma\in(0,\sigma_{*})$:
    ${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{\rho=0}^{\sigma}\mathcal{C}_{{\mathcal{R}}}(\rho;\widetilde{u}_{\rho})\;\!\mathrm{d}\rho={\mathcal{E}}(u^{\circ})\,.$ (3.17)
Later on, in Section 4.3 we will provide a sufficient condition for the attainment of the infimum in the definition (3.13) of $\mathcal{C}_{{\mathcal{R}}}$, which is also related to the validity of the analog of the lower semicontinuity properties (3.11b). This will lead to the existence of measurable selections
$(0,\infty)\ni\sigma\mapsto\widetilde{\xi}_{\sigma}\in\partial_{{\mathcal{R}}}{\mathcal{E}}(\sigma;\widetilde{u}_{\sigma}),\quad\text{with $\sigma\mapsto\widetilde{u}_{\sigma}\in\bm{J}(\sigma)$ a measurable selection,}$
@@
${\mathcal{E}}(\widetilde{u}_{\sigma})+\sigma{\mathcal{R}}\big(\frac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)+\int_{0}^{\sigma}{\mathcal{R}}^{*}({-}\widetilde{\xi}_{\rho})\;\!\mathrm{d}\rho\leq{\mathcal{E}}(u^{\circ})\,$,
@@
We conclude this section by using Example 2.14, revisited in the Banach setup, to convey the idea that the finer information encoded in the conditioned $\mathcal{R}$}-slope of the energy may play a key role for the attainment of the De Giorgi identity.
@@
We now treat Example 2.14 as a $\mathrm{gBGS}$ in the Hilbert or Banach space $X={\mathbb{R}}$:
$M={\mathbb{R}},\quad{\mathcal{E}}(u)=\max\{u,0\}=:u^{+},\quad{\mathcal{R}}(v)=\frac{1}{2}v^{2}.$
Since $\mathcal{E}$} is convex, its Fréchet subdifferential $\partial{\mathcal{E}}:{\mathbb{R}}\rightrightarrows$$\mathbb{R}$} coincides with the subdifferential in the sense of convex analysis, and it is set-valued with $\partial{\mathcal{E}}(0)=[0,1]$. The $\mathcal{R}$}-slope of the energy is given by
@@
As before, starting from $u^{\circ}=1$ we calculate the variational interpolant $\widetilde{u}_{\sigma}=(1{-}\sigma)^{+}=\max\{1{-}\sigma,0\}$, so that
@@
${\mathcal{E}}(u^{\circ})-{\mathcal{E}}(\widetilde{u}_{\sigma})-\sigma{\mathcal{R}}\big(\tfrac{1}{\sigma}(\widetilde{u}_{\sigma}{-}u^{\circ})\big)=1{-}(1{-}\sigma)^{+}-\tfrac{1}{2\sigma}|(1{-}\sigma)^{+}{-}1|^{2}=\begin{cases}\frac{1}{2}\sigma&\text{if }\sigma\leq 1,\\ 1{-}\frac{1}{2\sigma}&\text{if }\sigma\geq 1.\end{cases}$